in 2010, there were 50 fish in a pond. one year later, the number of fish decreased by 20%. how many fish were in the pond in 2011?
Answer by amarjeeth123(569)

The number of fish in 2011=50(1-20/100)=50(80/100)=40
The number of fish in 2011=40.